# The NP-hardness of the Gromov-Wasserstein distance

### v3

@@ Abstract
This note addresses the property frequently mentioned in the literature that the Gromov-Wasserstein (GW) distance is NP-hard. We provide the details on the non-convex nature of the GW optimization problem that imply NP-hardness of the GW distance between finite spaces for any instance of an input data. We further illustrate the non-convexity of the problem with several explicit examples.
@@
In order to perform GW based object comparison, one needs to compute a GW distance, a constrained minimization problem frequently mentioned as being NP-hard due to its non-convexity (see, e.g., discussion in Section 7 of [22], Section 1.1 of [25], Section 1 of [29], and/or Section 3 of [8]). To solve this problem in practice, two main approaches are currently taken in the literature, none of which guarantees a global minimum. The first approach is to employ the conditional gradient method (also known as Frank-Wolfe), which converges to a stationary point of the objective function that may not be a global minimum in the non-convex case (see Sections 2.1 and 2.2.2 of [4]). Conditional gradient for GW distance is implemented in the Python toolbox [17] and used for variants of GW problem in [7, 2]. The second approach proposed in [25] is to regularize the objective by adding an entropy term (see equation (7) of [25] and discussion in Section 10.6.4 in [24]). Such regularization does not make the problem convex222and hence no global minima are guaranteed when solving the problem via projected gradient descent proposed in [25] - see Remark 3 of [25], but allows for efficient implementations (Python toolboxes [17] and [12]) making the regularization technique widely used for various GW related problems [34, 9, 20, 28].
@@
This work provides the details on the structure of the GW optimization problem for finite spaces333the case most commonly encountered in real life applications that permit establishing the non-convex quadratic nature and imply NP-hardness of the GW distance. We note that the value of a GW distance is analytically determined in several special cases of metric-measure spaces with infinitely many points: see Theorem 4.1 [13] for GW distance between Gaussians when restricting the problem to Gaussian plans, and Theorem 1 of [1] for GW distance between spheres. For finite spaces, [23] proposes a polynomial time algorithm to compute an ultrametric variant of the GW distance with exponent $p=\infty$ (see Section 5.1 of the reference). Nevertheless, as the current note highlights, computing the $p$-GW distance, $p\in[0,\infty)$, between two datasets with finitely many points remains an NP-hard problem for any instance of an input data.
@@
The remainder of this note is organized as follows: Definition 1 defines a GW distance as an optimization program, followed by Example 3 that illustrates the structure of the program and its non-convexity on a pair of small-size exemplar metric spaces. Theorem 4 (main result) shows that GW distance between any two metric-measure spaces with finitely many points is NP-hard to compute. The result of the theorem is applied in Observation 5 and illustrated on the real life data in Example 6. The datasets and codes to reproduce Example 6 are available at https://github.com/kravtsova2/GW_NPhard.
@@
Given two finite metric-measure spaces $\mathcal{X}:=(X,d_{X},\mu_{X})$ and $\mathcal{Y}:=(Y,d_{Y},\mu_{Y})$ with points $X=\{x_{1},\ldots,x_{m}\}$ and $\{y_{1},\ldots,y_{n}\}$ respectively, define $GW_{p}^{p}(\mathcal{X},\mathcal{Y})$, $p\in[1,\infty)$, by the optimal value of the optimization problem444see also equation (4) of [25] and/or equations (1) and (2) of [29] that use the same notation as we use here.
@@
(Spaces are from Example 5.1 of [22])). Consider the space $\mathcal{X}:=(X,d,\mu_{X})$ with two points $X:=\{x_{1},x_{2}\}$ with distance $d$ giving $(x_{1},x_{2})=d(x_{2},x_{1})=1$ and $d(x_{1},x_{1})=d(x_{2},x_{2})=0$ (this can be conveniently represented by a distance matrix $(d_{ik})=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$) and measure $\mu_{X}:=(\frac{1}{2},\frac{1}{2})$. A pictorial representation of such space is provided in Figure 1A, top panel. Similarly, consider the space $\mathcal{Y}:=(Y,h,\mu_{Y})$ with points $Y:=(y_{1},y_{2})$, distance $(h_{jl})=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and measure $\mu_{Y}:=(\frac{1}{4},\frac{3}{4})$.
@@
Theorem 4 below shows that the presence of negative eigenvalues in the GW objective matrix is not specific to Example 3, but rather holds in general when posing GW problem between any pair of metric-measure spaces with finitely many points. The result states that the matrix in the objective function of GW distance from Definition 1 cannot be positive semidefinite, regardless of an input data. Hence, the resulting quadratic program is not convex and thus is NP-hard from computational perspective ([30]; for a recent review on the literature on computational complexity results for quadratic programming, see [15]).
Theorem 4 (GW distance between finite spaces is NP-hard).
The GW distance program in Definition 1 is non-convex quadratic (and hence is NP-hard).
@@
By Sylvester’s criterion stating that all principal minors of a real symmetric positive semidefinite matrix must be non-negative (see, e.g., [26]), the matrix $\Gamma$ or, more generally, $\Gamma_{p}$, - cannot be positive semidefinite. By the discussion preceding the theorem, a quadratic program with such objective cannot be convex and hence is NP-hard. ∎
Theorem 4 indicates that every instance of the GW distance optimization program is NP-hard, which supports the ongoing efforts on its computation (see, e.g., [23] for a recent review). At the same time, GW distance can itself be viewed as a relaxation of the Gromov-Hausdorff distance problem (see Section 4.1 of [22]) that generalizes a Quadratic Assignment Problem (QAP) (see Remark 4.6 of [22] for the connection between Gromov-Hausdorff and QAP). Below we discuss the interpretation of GW as a relaxation of QAP in the form that QAP was introduced in the economic literature, and use Theorem 4 to argue that such relaxation is not convex.
Observation 5 (Relation to the Quadratic Assignment Problem).
@@
$=\underbrace{\sum_{ijkl}d_{ik}^{2}\cdot\mu_{ij}\mu_{kl}+\sum_{ijkl}h_{jl}^{2}\cdot\mu_{ij}\mu_{kl}}_{\text{depend on $\mu_{X}$ and $\mu_{Y}$ only, and thusare irrelevant for optimization}}-\sum_{ijkl}2d_{ik}h_{jl}\cdot\mu_{ij}\mu_{kl}$
$=\text{constant}+2\sum_{ijkl}\underbrace{(-d_{ik})}_{\text{call }f_{ik}}h_{jl}\cdot\mu_{ij}\mu_{kl}$
@@
Acknowledgements and Funding The author thanks Facundo Mémoli for suggesting this problem, providing guidance during the solution process, and supporting the author with NSF DMS #2301359 and NSF CCS #1839356. The author thanks Adriana Dawes for the advice and support. A version of the problem considered in this note was given as an assignment during the topics course Optimal Transport based methods in Data Science (Math8610) taught by Facundo Mémoli at The Ohio State University in the Spring 2023 semester (https://github.com/ndag/OT-DS).

### v4

# Note on computational complexity of the Gromov-Wasserstein distance

@@ Abstract
This note addresses computational difficulty of the Gromov-Wasserstein distance frequently mentioned in the literature. We provide details on the structure of the Gromov-Wasserstein optimization problem that show its non-convex quadratic nature for any instance of an input data. We further illustrate the non-convexity of the problem with several explicit examples.
@@
In order to perform GW based object comparison, one needs to compute a GW distance, a constrained minimization problem frequently mentioned to be computationally difficult due to its non-convexity (see, e.g., discussion in Section 7 of [21], Section 1.1 of [27], Section 1 of [31], and/or Section 3 of [8]). This note provides the details on the structure of the GW optimization problem for finite spaces - the case most commonly encountered in real life applications - that demonstrate the non-convex quadratic nature of the GW distance on finite spaces.
A general instance of non-convex quadratic programming is NP-hard [32]. Furthermore, restrictions to several specific subclasses of non-convex quadratic programs were proved to be NP-hard as well. Among these are non-convex quadratic knapsack problem [25] (NP-hardness established using maximum clique problem222given a graph $G$, find a maximal size subset of vertices which are all connected) and quadratic problem with a single negative eigenvalue [24] (NP-hardness established using clique problem333given a graph $G$ and a positive integer $k$, decide whether $G$ has a subset of size $k$ of vertices that are all connected.). The former has specific constraints (the optimization variable is constrained to be a probability distribution), while the latter has specific objective structure (a single negative eigenvalue). At the same time, there are subclasses of non-convex quadratic programs that are shown to be solvable in polynomial time, e.g. a program with non-negativity constraints and a fixed (in terms of the length of optimization variable) number of negative eigenvalues [23]. Non-convex quadratic nature of GW distance that we illustrate here suggests that this subclass of non-convex quadratic programs could be NP-hard in its general instance. We are not aware of the proof that would utilize the reduction from a known NP-hard problem to establish NP-hardness of the GW distance similarly to establishing NP-hardness of the subclasses mentioned above.
@@
To solve the GW problem in practice, two main approaches are currently taken in the literature, none of which guarantees a global minimum. The first approach is to employ the conditional gradient method (also known as Frank-Wolfe), which converges to a stationary point of the objective function that may not be a global minimum in the non-convex case (see Sections 2.1 and 2.2.2 of [4]). Conditional gradient for GW distance is implemented in the Python toolbox [16] and used for variants of GW problem in [7, 2]. The second approach proposed in [27] is to regularize the objective by adding an entropy term (see equation (7) of [27] and discussion in Section 10.6.4 in [26]). Such regularization does not make the problem convex444and hence no global minima are guaranteed when solving the problem via projected gradient descent proposed in [27] - see Remark 3 of [27], but allows for efficient implementations (Python toolboxes [16] and [12]) making the regularization technique widely used for various GW related problems [36, 9, 19, 30].
@@
We note that the value of a GW distance is analytically determined in several special cases of metric-measure spaces with infinitely many points: see Theorem 4.1 [13] for GW distance between Gaussians when restricting the problem to Gaussian plans, and Theorem 1 of [1] for GW distance between spheres. For finite spaces, [22] proposes a polynomial time algorithm to compute an ultrametric variant of the GW distance with exponent $p=\infty$ (see Section 5.1 of the reference). We are not aware of a polynomial time algorithm that would solve a general instance of the $p$-GW distance, $p\in[0,\infty)$, between two datasets with finitely many points.
@@
The remainder of this note is organized as follows: Definition 1 defines a GW distance as an optimization program, followed by Example 3 that illustrates the structure of the program and its non-convexity on a pair of small-size exemplar metric spaces. Theorem 4 (main result) shows that GW distance between any two metric-measure spaces with finitely many points is a non-convex quadratic program. The result of the theorem is applied in Remark 5 and illustrated on the real life data in Example 6. The datasets and codes to reproduce Example 6 are available at https://github.com/kravtsova2/GW_computational.
@@
Given two finite metric-measure spaces $\mathcal{X}:=(X,d_{X},\mu_{X})$ and $\mathcal{Y}:=(Y,d_{Y},\mu_{Y})$ with points $X=\{x_{1},\ldots,x_{m}\}$ and $\{y_{1},\ldots,y_{n}\}$ respectively, define $GW_{p}^{p}(\mathcal{X},\mathcal{Y})$, $p\in[1,\infty)$, by the optimal value of the optimization problem555see also equation (4) of [27] and/or equations (1) and (2) of [31] that use the same notation as we use here.
@@
[…]
@@
Theorem 4 below shows that the presence of negative eigenvalues in the GW objective matrix is not specific to Example 3, but rather holds in general when posing GW problem between any pair of metric-measure spaces with finitely many points. The result states that the matrix in the objective function of GW distance from Definition 1 cannot be positive semidefinite, regardless of an input data. Hence, the resulting quadratic program belongs to a class of non-convex quadratic programs [32].
Theorem 4 (Non-convex quadratic nature of GW distance between finite spaces).
The GW distance program in Definition 1 is a non-convex quadratic optimization program.
@@
By Sylvester’s criterion stating that all principal minors of a real symmetric positive semidefinite matrix must be non-negative (see, e.g., [28]), the matrix $\Gamma$ - or, more generally, $\Gamma_{p}$, - cannot be positive semidefinite. By the discussion preceding the theorem, a quadratic program with such objective cannot be convex. ∎
Theorem 4 indicates that every instance of the GW distance optimization program is non-convex quadratic, which supports the ongoing efforts on its computation (see, e.g., [22] for a recent review). At the same time, GW distance can itself be viewed as a relaxation of the Gromov-Hausdorff distance problem (see Section 4.1 of [21]) that generalizes a Quadratic Assignment Problem (QAP) (see Remark 4.6 of [21] for the connection between Gromov-Hausdorff and QAP). Below we discuss the interpretation of GW as a relaxation of QAP in the form that QAP was introduced in the economic literature, and use Theorem 4 to argue that such relaxation is not convex.
Remark 5 (Relation to the Quadratic Assignment Problem).
@@
[…]
$\text{constant}+2\sum_{ijkl}\underbrace{(-d_{ik})}_{\text{call }f_{ik}}h_{jl}\cdot\mu_{ij}\mu_{kl}$
@@
Acknowledgements and Funding. The author thanks anonymous referees for pointing out deficiencies in the earlier version of the note. The author thanks Facundo Mémoli for suggesting this problem and supporting the author with NSF DMS #2301359 and NSF CCS #1839356. The author thanks Adriana Dawes for the advice and support. A version of the problem considered in this note was given as an assignment during the topics course Optimal Transport based methods in Data Science (Math8610) taught by Facundo Mémoli at The Ohio State University in the Spring 2023 semester (https://github.com/ndag/OT-DS). The author is solely responsible for all mistakes.